# CrashFixer: A crash resolution agent for the Linux kernel

### v1

@@ 1. Introduction
We have deployed CrashFixer and kGymSuite to Google Cloud and conducted over 500,000 kernel patching experiments. We evaluate CrashFixer’s capabilities on both kBenchSyz —a benchmark of curated Syzkaller bugs (Mathai et al., [n. d.])—and recent open Syzkaller bugs. On kBenchSyz, CrashFixer resolves $\sim 49\%$ of kernel crashes automatically with file-level localization assistance (i.e., when the file containing the developer fix is disclosed to CrashFixer). Providing a minimal, relevant execution trace preceding the crash, results in a relative performance increase of $\sim 10\%$. In an ongoing study of applying CrashFixer to resolve the latest unaddressed Syzkaller bugs , we find two plausible patches and continue to investigate other plausible solutions. A detailed manual analysis of the CrashFixer patches for $68$ kBenchSyz bugs shows that $(1)$ for $10$ bugs, CrashFixer generates accurate natural language diagnoses and nearly identical patches to the ground-truth developers’ fixes; $(2)$ for $9$ bugs, the generated diagnoses and patches effectively help pinpoint root causes; and $(3)$ for the remaining $49$ bugs, CrashFixer successfully (but perhaps questionably) resolves the crashes, but through “amputation” (i.e. bypassing faulty code all together) or through significant modification of existing functionality. This raises an interesting research question: how to guide LLMs to preserve intended defined behaviors of code when resolving undefined behaviors such as crashes, which we leave for future work.
@@ 6.1. Evaluation Setup
Models. Our methodology is LLM-agnostic. For cost efficiency, we use Gemini 1.5 Pro (versions 001 and 002) as the underlying LLM. The results reported may not reflect the performance of newer Gemini models or LLMs from other providers.
@@
2. Edit-Compile-Test 1 4 2 34.00 (32/36)
3. Edit-Compile-Test ++ 1 4 2 42.00 (42/42)
@@
10. Hyp-SR then Patch-SC $\mathcal{F}$** 4 4 1 75.00 (72/79)
Table 2. Average (min/max) number of crashes resolved in $3$ runs (SR: Self-reflection; SC: Self-consistency; $\mathcal{F}$: debug forests instead of debug tree; ** indicates Gemini-1.5-pro-002).
@@
Table 2 (row $7$) shows that CrashFixer resolves $\sim 58$ bugs of SCS crashes—5.8$\times$ more than the simple agent in kGym. It also resolves $\sim 65$ bugs of LCS crashes (row $9$). With a stronger model (Gemini 1.5-pro-002 rows $8$ and $10$), mitigated crashes increase to $\sim 62$ (SCS) and $\sim 75$ (LCS). Table 2 ablates different CrashFixer settings as described below.
@@
  • Introducing Forests (Table 2: rows 6–10). We see clear gains by decomposing the original tree into a forest of smaller, parallel $\mathbb{T}_{debug}$. Applying the Hyp then Patch strategy with 4 trees ($\mathcal{D}=4,\mathcal{B}=1$) raises performance to $\sim 54$ resolved bugs. This improvement stems from giving CrashFixer multiple fresh starting points, avoiding early edits that negatively impact downstream nodes. Adding self-consistency to select the patch best aligned with the generated hypothesis further boosts performance to $\sim 58$. This trend continues for the better model Gemini-1.5-pro-002 $\sim 62$ as well as long context bugs as well, up to $\sim 75$.
Adding Execution Traces. Only $255$ out of $279$ kBenchSyz bugs come with execution traces due to the reasons mentioned in Section 5.1.2. Thus, we evaluate the effectiveness of adding execution traces only on the $255$ bugs. In LCS, CrashFixer has a relative improvement of $10.36\%$ using 1.5-pro-002 and $13.61\%$ using 1.5-pro-001. In SCS, we see a relative increase of $9.7\%$ with 1.5-pro-002 and consistent results for 1.5-pro-001 (see Appendix A.2 for the detailed numbers).
Multi-turn. To understand the importance of multiple edit attempts, we measure the % of bugs only solved at depths greater than $1$ in SCS. For 1.5-pro-001, this accounts for $\sim 50\%$ of solved bugs in our final strategy (without execution) and $\sim 37.56\%$ with execution, demonstrating that tree-based exploration is broadly effective within the benchmark.
@@
Many plausible solutions can resolve a kernel crash, so predicting a patch equivalent to a developer’s is challenging. Sadly, a generated edit may resolve the kernel crash while compromising functionality. Thus, we manually analyze the patches for $68$ bugs in one run of Gemini 1.5-pro-002 (with execution) to understand the utility of the CrashFixer patches compared to the ground-truth fix patches for them.
Table 3 shows the utility results. We find that the patches fall into the following three categories. $(1)$ We consider patches as plausible if they are semantically equivalent to the developer patch. $10$ out of $68$ CrashFixer patches are plausible. $(2)$ We consider a patch as helpful if it targets the root cause correctly, but is not perfectly equivalent. $9$ CrashFixer patches are helpful. $(3)$ An incorrect patch resolves the crash but in the process modifies (or amputates) functionality. $49$ CrashFixer patches are incorrect, raising an interesting open research question of how to enable agents to verify that intended functionality is preserved after a crash is mitigated, in the absence of a practice of fine-grained unit tests in the kernel. We show examples of these categories in Section 7.
@@
68 10 9 49
Table 3. Patch Utility using 1.5-pro-002 (with execution).
@@
For a single kBenchSyz bug, with $4$ trees (having $4$ nodes each), we complete a maximum of $16$ debug cycles. Assuming the localized file content is $100K$ tokens, and that CrashFixer generates $8K$ tokens in each step of the pipeline, we estimate the average cost to be \$$\sim\$13$ for each bug. This estimate assumes that the LLM is always successful at getting a compilable patch without having to repeat the hypothesis-to-patch decomposition. In practice, this isn’t true and will vary for each LLM. Each debug cycle can take a maximum of $\sim 15$ minutes. However, as each tree can run in parallel, the average bug should ideally take $\sim 1$ hour to complete in the worst case, given enough machines.
@@ A.2. Evaluation Result with Execution Traces
In Table 4, we report our results when running CrashFixer with execution information. For a fair comparison, we report numbers on the subset of data for which our execution collection process works ($255$/$279$ data points). This “execution” subset removes $6$ SCS bugs and $18$ LCS bugs. As shown, when using execution traces, we see relative improvements in both LCS and SCS. In LCS, we see relative improvements of $10.36\%$ for 1.5-pro-002 and $13.61\%$ for 1.5-pro-001. In SCS, we see a $9.7\%$ relative increase for 1.5-pro-002 and consistent results for 1.5-pro-001.

### v2

@@ 1. Introduction
We have deployed CrashFixer and kGymSuite to Google Cloud and conducted over 500,000 kernel patching experiments. We evaluate CrashFixer’s capabilities on both kBenchSyz —a benchmark of curated Syzkaller bugs (Mathai et al., [n. d.])—and recent open Syzkaller bugs. On kBenchSyz, CrashFixer resolves $\sim 63\%$ of kernel crashes automatically with file-level localization assistance (i.e., when the file containing the developer fix is disclosed to CrashFixer). Providing a minimal, relevant execution trace preceding the crash, results in a relative performance increase of $\sim 9\%$. In an ongoing study of applying CrashFixer to resolve the latest unaddressed Syzkaller bugs , we find two plausible patches and continue to investigate other plausible solutions. A detailed manual analysis of the CrashFixer patches for $79$ kBenchSyz bugs shows that $(1)$ for $26$ bugs, CrashFixer generates accurate natural language diagnoses and nearly identical patches to the ground-truth developers’ fixes; $(2)$ for $12$ bugs, the generated diagnoses and patches effectively help pinpoint root causes; and $(3)$ for the remaining $41$ bugs, CrashFixer successfully (but perhaps questionably) resolves the crashes, but through “amputation” (i.e. bypassing faulty code all together) or through significant modification of existing functionality. This raises an interesting research question: how to guide LLMs to preserve intended defined behaviors of code when resolving undefined behaviors such as crashes, which we leave for future work.
@@ 6.1. Evaluation Setup
Models. Our methodology is LLM-agnostic. For cost efficiency, we use Gemini 2.5 Pro and 1.5 Pro (versions 001 and 002) as the underlying LLM. The results reported may not reflect the performance of newer Gemini models or LLMs from other providers.
@@
2. Edit-Compile-Test 1 4 2 34 (32/36)
3. Edit-Compile-Test ++ 1 4 2 42 (42/42)
@@
11. Hyp-SR then Patch-SC $\mathcal{F}$** 4 4 1 75 (72/79)
12. Hyp-SR then Patch-SC $\mathcal{F}$## 4 4 1 101 (100/102)
Table 2. Average (min/max) number of crashes resolved in $3$ runs (SR: Self-reflection; SC: Self-consistency; $\mathcal{F}$: debug forests instead of debug tree; ** indicates Gemini-1.5-pro-002; ## indicates Gemini-2.5-pro).
@@ 12. Hyp-SR then Patch-SC $\mathcal{F}$## 4 4 1 101 (100/102)
Table 2 (row $7$) shows that CrashFixer resolves $\sim 58$ bugs of SCS crashes—5.8$\times$ more than the simple agent in kGym. It also resolves $\sim 65$ bugs of LCS crashes (row $10$). With progressively stronger models (Gemini 1.5-pro-002 rows $8$ $11$ and Gemini 2.5 pro rows $9$ $12$), mitigated crashes increase steadily to $\sim 76$ (SCS) and $\sim 101$ (LCS). Table 2 ablates different CrashFixer settings as described below.
@@
  • Introducing Forests (Table 2: rows 6–12). We see clear gains by decomposing the original tree into a forest of smaller, parallel $\mathbb{T}_{debug}$. Applying the Hyp then Patch strategy with 4 trees ($\mathcal{D}=4,\mathcal{B}=1$) raises performance to $\sim 54$ resolved bugs. This improvement stems from giving CrashFixer multiple fresh starting points, avoiding early edits that negatively impact downstream nodes. Adding self-consistency to select the patch best aligned with the generated hypothesis further boosts performance to $\sim 58$. This trend continues for the better model Gemini-1.5-pro-002 $\sim 62$ and finally the best model Gemini-2.5-pro $\sim 76$.
Adding Execution Traces. We extract execution traces for $255$ out of $279$ kBenchSyz bugs due to the reasons mentioned in Section 5.1.2. Thus, we evaluate the effectiveness of adding execution traces only on these $255$ bugs. In LCS, CrashFixer has relative improvements of $14.6\%$ using 2.5-pro, $10.36\%$ using 1.5-pro-002, and $13.61\%$ using 1.5-pro-001. In SCS, we see a relative increase of $9.7\%$ with 1.5-pro-002 and consistent results for 1.5-pro-001 and 2.5-pro (see Appendix A.2 for the detailed numbers).
Multi-turn. To understand the importance of multiple edit attempts, we measure the % of bugs only solved at depths greater than $1$ in SCS. For 1.5-pro-001, this accounts for $\sim 50\%$ of solved bugs in our final strategy (without execution) and $\sim 37.56\%$ with execution, demonstrating that tree-based exploration is broadly effective within the benchmark. The corresponding numbers for 1.5-pro-002 are $35\%$ $27\%$ and for 2.5-pro are $33\%$ $24\%$ respectively.
@@
Many plausible solutions can resolve a kernel crash, so predicting a patch equivalent to a developer’s is challenging. However, a generated edit may resolve the kernel crash while compromising functionality. Thus, we manually analyze the patches for $79$ bugs in one run of Gemini 2.5-pro on the SCS (with execution) to understand the utility of the CrashFixer patches compared to the ground-truth fix patches for them.
Table 3 shows the utility results. We find that the patches fall into the following three categories. $(1)$ We consider patches as plausible if they are semantically equivalent to the developer patch. $26$ out of $79$ CrashFixer patches are plausible. $(2)$ We consider a patch as helpful if it targets the root cause correctly, but is not perfectly equivalent. $12$ CrashFixer patches are helpful. $(3)$ An incorrect patch resolves the crash but in the process modifies (or amputates) functionality. $41$ CrashFixer patches are incorrect, raising an interesting open research question of how to enable agents to verify that intended functionality is preserved after a crash is mitigated, in the absence of a practice of fine-grained unit tests in the kernel. We show examples of these categories in Section 7.
@@
79 26 12 41
Table 3. Patch Utility using 2.5-pro (with execution).
@@
For a single kBenchSyz bug, with $4$ trees (having $4$ nodes each), we complete a maximum of $16$ debug cycles. Assuming the localized file content is $100K$ tokens, and that CrashFixer generates $8K$ tokens in each step of the pipeline, we estimate the average cost to be \$$\sim\$13$ for each bug. This estimate assumes that the LLM is always successful at getting a compilable patch without having to repeat the hypothesis-to-patch decomposition. In practice, this isn’t true and will vary for each LLM. For 2.5-pro, we spend $\sim 21.62\$$ per bug (on average) in the SCS, after accounting for such retries.
Each debug cycle can take a maximum of $\sim 15$ minutes. However, as each tree can run in parallel, the average bug should ideally take $\sim 1$ hour to complete in the worst case, given enough machines.
@@ A.1. Details of Kernel Builds on kGymSuite
2.5-pro Hyp-SR then Patch-SC $\mathcal{F}$ 4 4 1 73.33 89
2.5-pro Execution + Hyp-SR/Patch-SC $\mathcal{F}$ 4 4 1 75 (72/79) 102 (102/102)
@@ A.2. Evaluation Result with Execution Traces
In Table 4, we report our results when running CrashFixer with execution information. For a fair comparison, we report numbers on the subset of data for which our execution collection process works ($255$/$279$ data points). This “execution” subset removes $6$ SCS bugs and $18$ LCS bugs. As shown, when using execution traces, we see relative improvements in both LCS and SCS. In LCS, we see relative improvements of $14.6\%$ for 2.5-pro, $10.36\%$ for 1.5-pro-002, and $13.61\%$ for 1.5-pro-001. In SCS, we see a $9.7\%$ relative increase for 1.5-pro-002 and consistent results for 1.5-pro-001 and 2.5-pro.